OpenStudy (khally92):

The force between two charges Q and q is 0.000000031N. when the two charges are separated by 3.0m. using the formula F=kQq/r^2 determine the force (F) when the separation (R) between two charges Q and q is increased from 3m to 4.5m.

OpenStudy (khally92):

its under variation

OpenStudy (abb0t):

since the distance changes from 3 to 4.5 \[\frac{ 4.5 }{ 3 } = 1.5\] and \[(1.5)^2 \] since it's: \[r^2\] Meaning the force is decrease by a factor of 2.25 Hence, \[\frac{ 3.1 \times 10^{-8} }{ 2.25 } = \]